OpenStudy (anonymous):

To which graph does the point (−1, 3) belong? y ≥ 2x + 6 y ≥ 3x − 1 y ≥ −x + 2 y ≥ −2x + 5

OpenStudy (anonymous):

plug your numbers into the inequalities and see which inequality becomes true.

OpenStudy (anonymous):

I will do the first one to show you... (-1,3) x = -1, y = 3 y >= 2x + 6 3 >= 2(-1) + 6 3 >= -2 + 6 3 >= 4 This is not true. So this one is not your solution. And just do the others this way to find which one is your solution

OpenStudy (anonymous):

@Kelliegirl33 i got it thanks . :)

OpenStudy (anonymous):

good to hear :)